A solution contains 4 grams of sodium hydroxide in 250ml of solution. If Na=23, O=16 , H=1 and
1cm3
of water weighs1g, what is the concentration of the solution in weight per weight percent?
1
Expert's answer
2015-10-21T06:00:38-0400
w = m(NaOH)/m(solution) *100% = 4/(250+4) = 1.57%

Answer: 1.57%